### v1

# Deriving pulsar pair-production multiplicities from pulsar wind nebulae using H.E.S.S. and LHAASO observations

Abstract
Context. Pulsar Wind Nebulae (PWNe) dominate the galactic gamma-ray sky at very high energies, and are major contributors to the leptonic cosmic ray flux. However, whether or not pulsars also accelerate ions to comparable energies is not yet experimentally confirmed.
Aims. We aim to constrain the birth period and pair-production multiplicity for a set of pulsars. In doing so, we aim to constrain the proportion of ions in the pulsar magnetosphere and hence the proportion of ions that could enter the pulsar wind.
Methods. We estimate possible ranges of the value of the average pair production multiplicity for a sample of 26 pulsars in the Australia Telescope National Facility (ATNF) catalogue, which have also been observed by the High Energy Stereoscopic System (H.E.S.S.) telescopes. We then derive lower limits for the pulsar birth periods and average pair production multiplicities for a subset of these sources where the extent of the pulsar wind nebula and surrounding supernova shell have been measured in the radio. We also derive curves for the average pair production multiplicities as a function of birth period for sources recently observed by the Large High Altitude Air Shower Observatory (LHAASO).
Results. We show that there is a potential for hadrons entering the pulsar wind for most of the H.E.S.S. and LHAASO sources we consider, dependent upon the efficiency of luminosity conversion into particles. We also present estimates of the pulsar birth period for six of these sources, which all fall into the range of $\simeq$10-50 ms.
Key Words.:
Pulsars: general, Pulsars: individual, Gamma rays: general
1 Introduction
Pulsar wind nebulae (PWNe) occur when, following a supernova explosion, a pulsar’s ultra-relativistic wind flows into the surrounding medium producing a wind termination shock that can accelerate particles to relativistic energies (Gaensler & Slane, 2006; Kargaltsev et al., 2015). PWNe are the most numerous class of galactic very-high-energy gamma-ray emitters (representing around $\sim$40of known galactic sources), and have been observed across the electromagnetic spectrum for $\simeq$50 years (Wakely & Horan, 2008; H.E.S.S. Collaboration, 2018b). A key question regarding PWNe is whether the gamma-ray emission produced is solely a result of inverse Compton scattering of relativistic electrons on background fields, or if photopion production caused by protons which originated from the pulsar could contribute to the emission at the highest energies (Bednarek & Protheroe, 1997; Bednarek, 2003).
A key metric which governs whether hadrons can likely escape the pulsar surface into the PWN is the average pair production multiplicity $\langle\kappa\rangle$. This describes the number of e+/e- pairs that escape the pulsar light cylinder per electron that escapes from the pulsar’s surface (as a result of cascades produced by gamma-ray Bremsstrahlung). This places some constraints on potential hadrons originating from the pulsar, as such particles do not multiply in cascades in the pulsar magnetosphere, and so can only make up a fraction of $1/\langle\kappa\rangle$ of the total particles at most (Kirk et al., 2007). Therefore for sources with a high pair-production multiplicity there exists the possibility of hadrons escaping into the wind. In this paper, we aim to determine the value of $\langle\kappa\rangle$ for a number of gamma-ray-emitting PWNe as observed by the High Energy Stereoscopic System (H.E.S.S.) gamma-ray telescopes (Aharonian, F. et al., 2006). This represents roughly a quarter of known PWN (Ferrand & Safi-Harb, 2012; Giacinti et al., 2020). We also make use of radio observations, as these can help to constrain the size of the PWN and surrounding Supernova Remnant (SNR) even once the associated gamma-ray emission has faded. Through this we can also place constraints on the pulsar birth period for a subset of these systems. We will also place constraints on the potential values of $\langle\kappa\rangle$ for a number of sources recently observed by the Large High Altitude Air Shower Observatory (LHAASO) (Cao et al., 2022). This paper builds upon the previous work by de Jager (2007), who introduced the concept of deriving pair-production multiplicity constraints from gamma-ray data and comparing them to estimates of the pulsar birth period obtained from radio data, but we utilise a greater number of considered sources, updated observations and updated modelling.
The remainder of this paper is organised as follows: in Section 2 we introduce the theoretical background to our models and discuss the input parameters to them, in Section 3 we present our main results. In Section 4 we discuss the implications for cosmic ray production, the effect of varying free parameters in the models and the prospects for future observations. Finally, in Section 5 we conclude.
2 Modelling theory
2.1 Deriving $P_{0}$
Following the reasoning of van der Swaluw & Wu (2001), the initial spin period of a pulsar embedded in a PWN inside a larger SNR can be determined for systems aged $\sim 10^{3}-10^{4}$ yr using the relation
$P_{0}=2\pi\left[\frac{2E_{0}}{\eta_{1}I}\left(\frac{R_{PWN}}{\eta_{3}R_{SNR}}\right)^{3}+\left(\frac{2\pi}{P_{t}}\right)^{2}\right]^{-1/2},$ (1)
where $E_{0}$ is the total mechanical energy of the Supernova Remnant (SNR) which we take to be $10^{51}\,\mathrm{erg}$, $I\approx 1.4\times 10^{45}\,\mathrm{g\,cm^{2}}$ is the moment of inertia of the neutron star (Mølnvik & Østgaard, 1985), $R_{SNR}$ and $R_{PWN}$ are the SNR and PWN radii as measured in the radio (as these mark the ‘true’ extent of the PWN (van der Swaluw & Wu, 2001)) and $P_{t}$ is the period of the pulsar at the present time. $\eta_{1}$ and $\eta_{3}$ are dimensionless scaling parameters that relate the radius of the PWN $R_{PWN}$ to the ratio of the SNR $R_{SNR}$
$R_{PWN}(t)=\eta_{3}(t)(\eta_{1}E_{SD}/E_{0})^{1/3}R_{SNR}(t).$ (2)
where $E_{SD}$ is the total spin-down energy of the pulsar. $\eta_{1}$ effectively a proxy for the strength of synchrotron losses, neglecting these and setting $\eta_{1}=1$ and $\eta_{3}=1.02$ is suggested by van der Swaluw & Wu (2001), leading to a systematic underestimation of the initial spin rate which propagates through to our estimates of $\langle\kappa\rangle$. Equations 1 and 2 are only valid in the subsonic phase of the PWN’s evolution, after most of the energy in the pulsar wind has been deposited into the PWN, and follow the assumption that the PWN/SNR system is spherically symmetric (van der Swaluw & Wu, 2001). de Jager (2007) also considers the work of van der Swaluw & Wu (2001), although they ultimately adopt a fixed $P_{0}$ in their multiplicity analysis as a constraint of radio observations available at the time. We utilise a $P_{0}$ value calculated for each H.E.S.S. source for which we now have radio data in our analysis.
2.2 Determining $N_{el}$ and $\langle\kappa\rangle$
The number of Goldreich-Julian electrons at the current time $t$, which represents the total number of electrons that have been stripped from the polar caps of the pulsar, is given by the integral
$N_{GJ}=\int_{t=0}^{t=-\tau(P_{0})}\frac{[6c\dot{E}(t)]^{1/2}}{e}(-dt)\,$ (3)
(Goldreich & Julian, 1969). To calculate $\dot{E}$, we take values of $\dot{E}(t)$ today as calculated in Giacinti et al. (2020), and assume that the energy output of the pulsars evolve as
$\dot{E}(t)=\dot{E}_{0}\left(1+\frac{t}{\tau_{0}}\right)^{-\alpha},$ (4)
with $\dot{E}_{t}$ the value of $\dot{E}$ at the current time, the constant $\tau_{0}$ set to $10^{3}\,$years and $\alpha=(n+1)/(n-1)$ assumed to be equal to 2 for a braking index $n=3$. $N_{GJ}$ is then integrated numerically. The average pair production multiplicity is then
$\langle\kappa\rangle=\frac{N_{el}}{2N_{GJ}}$ (5)
(de Jager, 2007).
We estimate the number of electrons in the PWNe observed by H.E.S.S. (Aharonian et al., 2005; H.E.S.S. Collaboration, 2018b) by following the approach of (Giacinti et al., 2020) and approximating the electron spectrum as a broken power law with a single spectral parameter (BPL1) such that
$N_{el}=E_{tot}\left(\frac{(2-\Gamma)E_{0}^{1-\Gamma}}{E_{2}^{2-\Gamma}-E_{1}^{2-\Gamma}}\right)\times\left(\frac{(E_{2}/E_{0})^{1-\Gamma}-(E_{1}/E_{0})^{1-\Gamma}}{1-\Gamma}\right)$ (6)
where $\Gamma=2.2$, $E_{0}$ is assumed to be $0.1\,\mathrm{TeV}$, $E_{1}$ is the energy threshold of the gamma-ray observations (taken either from the H.E.S.S. galactic plane survey (H.E.S.S. Collaboration, 2018a) or (Aharonian et al., 2005) for G0.9+0.1/PSR J1747-2809) and $E_{2}$ is assumed to be $10\,\mathrm{TeV}$. This is as the radiative lifetime of electrons above this energy is likely to be less than the typical ages of pulsars in our sample, since this scales with electron energy $E_{e}$ as $\sim 10^{4}(B/10\,\mathrm{\mu G})^{-2}(E_{e}/10\,\mathrm{TeV})^{-1}\,\mathrm{yr}$ (Giacinti et al., 2020). The values for the total energy in electrons $E_{tot}$ are taken from the modelling in Giacinti et al. (2020). We then calculate the age $\tau$ as a function of $P_{t}$ the pulsar period today, the period derivative today $\dot{P}_{t}$, the initial period $P_{0}$ and the braking index $n$ such that
$\tau(P_{t},\dot{P}_{t},P_{0},n)=\left(1-\left(\frac{P_{0}}{P_{t}}\right)^{n-1}\right)\times\frac{P_{t}}{(n-1)\dot{P}_{t}}$ (7)
(Gaensler & Slane, 2006).
Given the dependence of Equation 3 upon $\tau(P_{0})$ and therefore $P_{0}$ by Equation 7, we can find curves of $\langle\kappa\rangle$ as a function of $P_{0}$ (where $P_{0}$ is constrained to be shorter than the measured $P_{t}$ today). As these curves do not consider electrons producing emission outside the gamma-ray range, these curves serve as strict lower limits for the true pulsar multiplicity. This technique to determine these curves was first presented in (de Jager, 2007). As we can simultaneously find $P_{0}$ through the ratio of $R_{PWN}/R_{SNR}$ and Equation 1, for sources where $P_{t},R_{PWN},R_{SNR}$ are known and $N_{el}$ can be modelled, one can estimate a specific value for the lower bound of the pair production multiplicity. This is found by finding the intersection between the estimate for $P_{0}$ using Equation 2, and the curves of $\langle\kappa\rangle(P_{0})$ as estimated using Equations 3, 6 and 7111Jupyter notebooks showing this analysis can be found at www.github.com/STSpencer/psrprops..
We also derive pair-production multiplicity curves for those sources observed by LHAASO with published results of modelling of their gamma-ray emission. Note that these publications make different assumptions about, for example, the associated radiation fields compared to the H.E.S.S. sources. For the Dragonfly Nebula, associated with PSR J2021+3651, we utilise a power law model with an exponential cutoff (PLEC), such that
[…]
following Woo et al. (2023). Here, we use a total energy of $3.9\times 10^{48}\,\mathrm{erg}$ to determine the proportionality constant numerically, a reference energy of $E_{1}=\mathrm{1\,TeV}$, a lower energy threshold of $E_{0}=\mathrm{25\,TeV}$, a cutoff of $E_{cut}=\mathrm{900\,TeV}$ and a maximum energy of $E_{2}=1400\,\mathrm{TeV}$ following Cao et al. (2023) and a spectral index of $\Gamma=\mathrm{1.4}$. For PSR J1841-0345 we also follow Albert et al. (2023) in using a power law spectrum with an exponential cutoff, with a proportionality constant of $7.1\times 10^{31}\,\mathrm{erg}$, a low-energy threshold of $E_{0}=\mathrm{10\,MeV}$, a reference energy of $E_{1}=\mathrm{7\,TeV}$, a cutoff energy of $E_{cut}=\mathrm{72\,TeV}$, a maximum energy of $E_{2}=\mathrm{740\,TeV}$ and a spectral index of $\Gamma=\mathrm{2.2}$. For PSR J1849-0001 we follow Amenomori et al. (2023) with a power-law model such that
$N_{el}\propto\left[\frac{1}{1-\Gamma}\left(\frac{E_{2}}{E_{0}}\right)^{1-\Gamma}-\frac{1}{1-\Gamma}\left(\frac{E_{1}}{E_{0}}\right)^{1-\Gamma}\right]$ (9)
with a low energy threshold of $E_{0}=\mathrm{0.5\,TeV}$, a reference energy of $E_{1}=\mathrm{10\,TeV}$, and high energy threshold of $E_{2}=\mathrm{100\,TeV}$ a spectral index of $\Gamma=2.5$ and a total energy of $2.8\times 10^{47}\,\mathrm{erg}$ from which the proportionality constant is also derived numerically. For PSR J1826-1334 we follow H. E. S. S. Collaboration et al. (2019) in using a broken power-law model with a two spectral parameters (BPL2) such that
[…]
with a low-energy threshold of $E_{0}=\mathrm{0.7\,TeV}$ (H.E.S.S. Collaboration, 2018a), a break energy of $E_{1}=\mathrm{0.9\,TeV}$, a maximum energy of $E_{2}=\mathrm{42\,TeV}$ a lower-energy spectral index of $\Gamma=1.4$ and a high-energy spectral index of $\Gamma_{2}=3.25$ with a total energy of $5.5\times 10^{48}\,\mathrm{erg}$ from which the proportionality constant is determined numerically. A summary of the input parameters to our modelling, along with the spectral models used for each source, are shown in Table 2. We obtained values for $R_{SNR}$ and $R_{PWN}$ from (Giacinti et al., 2020) and references therein.
Table 1: Input parameters to our modelling and the corresponding spectral model used.
ATNF Name $R_{SNR}$ $\mathrm{[pc]}$ $R_{PWN}$ $\mathrm{[pc]}$ $P_{t}$ $\mathrm{[ms]}$ $\dot{P}_{t}$ [×$\mathrm{[\times 10^{-13}\ s/s]}$ Model $E_{0}$ $\mathrm{[TeV]}$ $E_{1}$ $\mathrm{[TeV]}$ $E_{cut}$ $\mathrm{[TeV]}$ $E_{2}$ $\mathrm{[TeV]}$ $\Gamma$ $\Gamma_{2}$ Refs.
J1833-1034 2.98 0.8 61.8 2.02 BPL1 0.1 0.39 - 10 2.2 - $1,2$
J1513-5908 38.4 19.2 151.6 15.3 BPL1 0.1 0.61 - 10 2.2 - $1,2$
J1930+1852 10.8 2.7 136.9 7.50 BPL1 0.1 0.89 - 10 2.2 - $1,2$
J1846-0258 2.6 0.58 326.6 71.1 BPL1 0.1 0.40 - 10 2.2 - $1,2$
J0835-4510 19.5 12.2 89.3 1.25 BPL1 0.1 0.61 - 10 2.2 - $1,2$
J1747-2809 19.8 2.5 52.2 1.56 BPL1 0.1 0.17 - 10 2.2 - $1,3$
J2021+3651 - - 103.7 0.957 PLEC 1 25 900 1400 1.4 - $4$
J1841-0345 - - 112.9 1.55 PLEC $1\times 10^{-5}$ 7.0 72 740 2.2 - $5$
J1849-0001 - - 38.5 0.142 PL 0.5 10 - 100 2.5 - $6$
J1826-1334 - - 101.5 0.753 BPL2 0.7 0.9 - 42 1.4 3.25 $7$
[…]
(1) Giacinti et al. (2020), (2) H.E.S.S. Collaboration (2018a), (3) Aharonian et al. (2005), (4) Woo et al. (2023), (5) Albert et al. (2023), (6) Amenomori et al. (2023), (7) H. E. S. S. Collaboration et al. (2019)
3 Results
For 26 pulsars in the Australia Telescope National Facility ATNF catalogue for which the region has been observed by H.E.S.S., we can place constraints on the lower limit of $\langle\kappa\rangle$. We do this by assuming that the birth period is in the range 10-50 ms, based on the reasoning that pulsars with birth periods below 10 ms are unlikely to exist (Perna et al., 2008), and PWNe associated with longer period pulsars are unlikely to produce gamma-ray emission. The resulting ranges for the possible lower-limit of $\langle\kappa\rangle$ are shown in Figure 1, with numerical values provided in Appendix A.
[…]
$E_{\mathrm{CR}}\approx 1.2\times 10^{20}A_{56}\eta\langle\kappa\rangle_{\rm lim,4}I_{45}B_{13}^{-1}R_{\star,6}^{-3}\tau_{7.5}^{-1}$ (11)
(Kotera et al., 2015), where $E_{\mathrm{CR}}$ is the iron nuclei energy which we assume to be $3\,\mathrm{PeV}$, roughly equivalent to the cosmic ray ‘knee’ (Blasi, 2013; Grenier et al., 2015), $A_{56}=1$ is the mass number of the stripped particles normalised to 56, and $B,R_{\star},I$ and $\tau$ are the magnetic field strength, radius of the pulsar, moment of inertia and age of the system respectively (with subscripts $x$ representing normalisation to $10^{x}$ in cgs units). $I_{45}$ and $R_{\star,6}$ we assume to be 1 as they are non-trivial to determine (Özel & Freire, 2016), $\tau_{7.5}$ we calculate using equation (7) and $B_{13}$ we obtain using the equation
$B=3.2\times 10^{19}\sqrt{P_{t}\dot{P_{t}}}\,\mathrm{G}$ (12)
(Gaensler & Slane, 2006). $\eta\leq 1$ in equation (11) is a luminosity conversion efficiency factor that is effectively unknown. Previous theoretical studies considering hadronic acceleration by pulsars have used values ranging from 0.1 to 1 (Kotera, 2014), and it could feasibly be of order 0.01 (de Oña Wilhelmi et al., 2022), although de Oña Wilhelmi et al. (2022) show that for many known gamma-ray sources it appears to be in the range 0.1 to 1 for particles downstream of the wind termination shock. As such, the conclusion reached for whether hadrons can escape into the pulsar wind for a given pulsar depends on the adopted value of $\eta$, as the intersection point between our derived $P_{0}$ value and the curve of $\langle\kappa\rangle(P_{0})$ can fall above or below the derived value for $\langle\kappa\rangle_{\rm lim}$.
The derived initial periods, pair multiplicities and values for $\langle\kappa\rangle_{\rm lim}$ for the six H.E.S.S. sources where we have both an estimate of $R_{PWN}/R_{SNR}$ and $N_{el}$ are shown in Figure 2 for assumed values for $\eta$ of 1 and 0.1. The calculated values for $P_{0}$ and $\langle\kappa\rangle$ for these sources is summarised in Table 3. We can confidently not exclude the possibility of hadronic escape into the wind for the pulsars J0835-4510 and J1930+1852, as they have intersection points between their $\langle\kappa\rangle(P_{0})$ curve and the modelled value of $P_{0}$ below the marker representing $\langle\kappa\rangle_{\rm lim}$ in the maximal case of assuming $\eta=1$. However, we also cannot completely exclude this scenario for the other four H.E.S.S. pulsars depending on the $\eta$ value assumed. That said, for J1747-2809 $\eta$ would be required to be order $10^{-3}$, so it is unlikely that hadrons escape into the wind for this source. The result that a higher efficiency factor for the production of hadrons at the pulsar corresponds to a greater likelihood of said hadrons escaping into the wind is expected, but our results quantify the effect of this for the first time. A previous claim in Kotera et al. (2015) that there is a rough value of $\langle\kappa\rangle_{\rm lim}$ of $2m_{p}/m_{e}\approx 3672$ for determining if hadrons can escape into the wind, for the specific hypothetical case of a young, $P_{0}\leq 10\,\mathrm{ms}$ pulsar producing ultra-high-energy cosmic rays, is probably therefore only correct to within 1-2 orders of magnitude. This is because, as can be seen from Figure 2, the constraints on hadronic escape reached for any particular source have that order uncertainty, depending on the measured or assumed values of $P$, $P_{0}$ and $\eta$. Despite differences in our modelling, we obtain order-of-magnitude similarity to the curve derived in de Jager (2007) for PSR B1509-58 (PSR J1513-5908), and whilst for the Vela Pulsar (PSR J0835-4510) our lower limit curve is within the permitted range predicted in that work, the precise level of agreement depends on scaling to the magnetic field strength in the region. However, some discrepancies for the results for these two pulsars considered in both our work and de Jager (2007) are to be expected, as we base our results on more recent observational data and PWN modelling, as referenced in Table 2.
With the exception of PSR J1833-1034, we obtain birth periods that are consistent within a factor two of previous estimates. The birth period we derive for PSR J1833-1034 of 33.0 ms is in stark contrast to the value of ¿55 ms derived by Camilo et al. (2006); this is likely due to Equation 1 not being suitable for a source that is potentially of age $<1$ kyr (much lower than its characteristic age of $\tau_{c}\approx 4.8$ kyr) as suggested by Camilo et al. (2009). The validity of our pair-production multiplicity for this source is therefore also questionable. The birth period for PSR J1513-5908 we derive (15.2 ms) is in good agreement with the value derived by Abdo et al. (2010) of 16 ms. Similarly for the Vela Pulsar PSR J0835-4510 the value of 10.9  ms we obtain is in reasonable agreement with the value from Helfand et al. (2001) of 6 ms. The birth period of 47.9 ms we obtain for PSR J1747-2809 is consistent with the prediction of Camilo et al. (2009) who proposed it to be ¿40 ms. It has been postulated that pulsars with approximately millisecond birth periods could explain the ultra-high energy all-particle cosmic ray flux if they exist (Kotera et al., 2015); we find no pulsars with a $\sim\mathrm{ms}$ birth period in our sample. This is consistent with some previous studies utilising archival X-ray data (Perna et al., 2008).
The results of the modelling of the four LHAASO sources can be seen in Figure 3. We cannot attempt to derive values for $\langle\kappa\rangle_{\rm lim}$ as we do not have radio observations for these targets. But if we assume the potential range of $\langle\kappa\rangle_{\rm lim}$ values for these systems is comparable to those we have determined with the H.E.S.S. data, we cannot exclude the possibility of hadronic particles reaching the pulsar wind for any of these systems either.
Figure 1: Maximum and minimum values of the lower limit for $\langle\kappa\rangle$ for the pulsars for which we can estimate $N_{el}$. The minimum bounds for this lower limit correspond to assuming a birth period of 10 ms, and the maximum corresponds to assuming a birth period of 50 ms.
Figure 2: The intersections of our estimates for $P_{0}$ using Equation 2, and the curves of $\langle\kappa\rangle(P_{0})$ as estimated using Equations 3, 6 and 7 (the latter is constrained such that $P_{0}<P_{t}$) are shown by circular markers. The intersection location represents the lower limit for the pair-production multiplicity, see Table 3 for the associated values. Derived values for the pair-production multiplicity limit $\langle\kappa\rangle_{\rm lim}$ for each pulsar are shown as lower limit markers in matching colours to the multiplicity curve. The top panel assumes a luminosity conversion efficiency of $\eta=1$ in the determination of $\langle\kappa\rangle_{\rm lim}$, whereas the lower panel assumes $\eta=0.1$.
Table 2: Derived values of $P_{0}$, along with $N_{el}$, the total energy in electrons $E_{\mathrm{e,tot}}$, and the lower limit for $\langle\kappa\rangle$ for the six H.E.S.S. sources where this is possible.
ATNF Name $P_{0}$ $\mathrm{[ms]}$ $\dot{E}_{t}$ [×$\mathrm{[\times 10^{36}\ erg/s]}$ $N_{el}$ [×$\mathrm{[\times 10^{47}\ Counts]}$ $E_{\mathrm{e,tot}}$ [×$\mathrm{[\times 10^{47}\ erg]}$ $\langle\kappa\rangle$ $\mathrm{[Dimensionless]}$
J1833-1034 33.0 $33.9$ $45.6$ $51.8$ 1476
J1513-5908 15.2 $17.0$ $5.14$ $8.35$ 809
J1930+1852 41.3 $12.0$ $5.06$ $11.0$ 432
J1846-0258 50.8 $8.13$ $1.62$ $1.87$ 1061
J0835-4510 10.9 $6.92$ $18.7$ $1.87$ 174
J1747-2809 47.9 $42.7$ $125$ $71.4$ 29328
[…]
Figure 3: The calculated lower limits on the pair-production multiplicity for the four pulsars co-incident with LHAASO sources.
4 Discussion
4.1 Implications for production of cosmic rays
We cannot safely exclude the possibility of hadrons escaping the pulsar surface into the wind for any of our considered sources. Though it should be noted that the values of $\langle\kappa\rangle$ we observe for all sources are within likely theoretical constraints (given $P_{0}$ is not likely to be $\gg$50ms), which cap the maximum pair production multiplicity to be a few $10^{5}$ (Timokhin & Harding, 2019). The $\langle\kappa\rangle$ curve we observe for PSR J1849-0001 is significantly below unity, but similar findings were encountered by de Jager (2007), who found that the Vela Pulsar (PSR J0835-4510) had a pair production multiplicity less than unity (depending on the assumptions made).
4.2 The phase-space of free parameters for the H.E.S.S. sources
As our work is based on the previous findings of Giacinti et al. (2020), which did not provide estimates of the uncertainties on the PWN and SNR radii, we cannot perform a full, robust uncertainty analysis. Instead, to investigate the effect of various assumptions in our models, we examine the effect of changing parameters in our model for the two pulsars PSR J1747-2809 and PSR J1846-0258 (to probe the minimum and maximum values of $P_{t}$ respectively). The results of this are shown in Figures 4 and 5.
The selected value of $I$ appears to have an approximately order of magnitude effect on the point of intersection between the pair production multiplicity curve and the pulsar birth period estimate for the short period pulsar PSR J1747-2809 seen in Figure 4, but very little influence upon the derived pair-production multiplicity as seen in Figure 5 for PSR J1846-0258. Variations in the moment of inertia $I$ correspond to differences in the internal structure, and can be used to constrain the equation of state for dense matter Ravenhall & Pethick (1994). The results for both pulsars suggest there is a roughly factor 2 impact on the derived $P_{0}$ within the band of reasonable potential values. Changing the value of $E_{2}$ is seen to only have a modest effect on the value of $\langle\kappa\rangle$ in Figures 4 and 5. Further increasing $E_{2}$ would lead to a resulting decrease in $\langle\kappa\rangle$. Choosing 10 TeV for this value appears to be reasonable as a lower limit given our relative agreement for the Vela pulsar and PSR B1509-58 with de Jager (2007), where the pair-production multiplicity is calculated differently. Changing the value of $\Gamma$ by 0.4 appears to have an approximately factor two effect on the derived $\langle\kappa\rangle$. The true pair production multiplicity value is very likely to be within this range, as a value of 2.2 is expected for particles at an ultra-relativistic shock undergoing isotropic scattering (Achterberg et al., 2001). A similar magnitude effect is seen when varying $n$ for both pulsars, this likely dominates the uncertainty for shorter birth period pulsars. But as this braking index has only been measured for a very small population of pulsars, by convention we adopt a value of 3 as has been adopted by much of the literature (Gaensler & Slane, 2006).
4.3 Prospects for future observations
The large beaming fraction (up to 0.92) for gamma-rays from high $\dot{E}_{t}$ pulsars (Johnston et al., 2020) suggests that observations of gamma-ray pulsars are nearly complete. This implies conclusions of this study regarding the contribution of pulsars to the hadronic cosmic ray spectra are unlikely to change with the advent of future gamma-ray instruments such as the Cherenkov Telescope Array (Cherenkov Telescope Array Consortium et al., 2019) or the Southern Wide-Field Gamma-Ray Observatory (Albert et al., 2019). That said, state-of-the-art and future radio observations by observatories such as MeerKAT (Goedhart et al., 2023) and the upcoming Square Kilometer Array (Carilli & Rawlings, 2004) could allow for more PWN and SNR radii, from systems with lower surface brightness and larger extent, to be measured. This means the pair production multiplicities and birth period could potentially be derived for more pulsars. This, in combination with multi-wavelength modelling, would allow for more stringent constraints on hadronic contributions to the gamma-ray emission from PWNe.
Figure 4: The effect of varying $I$ (top), $E_{2}$ (upper middle), $\Gamma$ (lower middle) and $n$ (bottom) for PSR J1747-2809, which has the lowest $P_{t}$ (47.9 ms) in our sample of pulsars with a constrained $P_{0}$.
Figure 5: The effect of varying $I$ (top), $E_{2}$ (upper middle), $\Gamma$ (lower middle) and $n$ (bottom) for PSR J1846-0258, which has the highest $P_{t}$ (326.6 ms) in our sample of pulsars with a constrained $P_{0}$.
5 Conclusions
In this study we have placed constraints on the pair production multiplicity for 26 pulsars observed by H.E.S.S., deriving lower limits for the average pair production multiplicities and birth periods for six of these pulsars. We also placed lower limits on the pair production multiplicities for four pulsars co-incident with sources observed by LHAASO based on their very-high-energy gamma-ray emission. We find that the possibility of hadrons escaping the pulsar into the wind cannot be conclusively excluded for any source we consider. Nevertheless, for J1747-2809 the pulsar for which we infer the highest pair-production multiplicity, the escape of hadrons into the wind is essentially ruled out for values of $\eta\geq 0.1$. This implies that hadrons are only released into the wind of J1747-2809 if $\eta\leq 0.1$, which is too low for any hadronic component to be energetically significant.
Acknowledgements.
This work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project Number 452934793.
Appendix A Numerical estimates of $\langle\kappa\rangle$
Numerical bounds for the pair-production multiplicities for the 26 pulsars in our combined H.E.S.S./ATNF sample are shown in Table 3.
Table 3: Estimates of the maximum and minimum value of $\langle\kappa\rangle$ for the H.E.S.S./ATNF sample.
ATNF Name $\dot{E}_{t}$ [×$\mathrm{[\times 10^{36}erg/s]}$ $\langle\kappa\rangle_{min}$ $\mathrm{[Dimensionless]}$ $\langle\kappa\rangle_{max}$ $\mathrm{[Dimensionless]}$
J1833-1034 33.9 884 4430
J1513-5908 17.0 802 938
J1930+1852 12.0 374 464
J1420-6048 10.0 130 495
J1846-0258 8.13 1030 1060
J0835-4510 6.92 173 333
J1838-0655 5.50 98 349
J1418-6058 4.90 107 159
J1357-6429 3.09 109 128
J1826-1334 2.82 72 119
J1119-6127 2.29 268 273
J1301-6305 1.70 64 73
J1747-2809 42.7 1880 59185
J1617-5055 15.8 316 1072
J1023-5746 11.0 304 435
J1856+0245 4.57 80 191
J1640-4631 4.37 351 386
J1709-4429 3.39 89 145
J1907+0602 2.82 57 90
J1016-5857 2.57 42 65
J1803-2137 2.19 85 111
J1809-1917 1.82 26 61
J1718-3825 1.29 13 39
J1028-5819 0.832 6 12
J1833-0827 0.575 5 12
J1857+0143 0.447 8 10
[…]